# Optimized neural forms for solving ordinary differential equations

### v1

@@ 6.1 Test suite
| 7 | 80 | 160 | 320 | 1000 | 72 | 144 | 288 |
6.2 Dataset generation
For each test problem, two sets of data points were generated, namely a training dataset of size that was used for training the neural forms, and a test dataset of size that was used for assessing the generalization quality. For the training dataset, two grid types were considered with data points :
1.
Equidistant points:
2.
Chebyshev points:
Each grid type provides training datasets with different point distributions, which can be beneficial for capturing the structure of the ODE’s solution. Besides the grid types above, Gauss-Legendre points were also tested, but they are omitted since yielded results almost identical to the Chebyshev case.
A suitable size of the training set is always problem dependent. Thus, it is typically determined through preliminary trial-and-error experimentation. Naturally, more complex problems require higher numbers of training points. Table 1 reports the corresponding numbers that were eventually selected in our experiments after the preliminary screening of the test problems. For each test problem, three levels (low, medium, and high) were considered for . On the other hand, the test set consisted of equidistant points in the interval for all test problems.
@@ 6.3 Neural network architecture
where is the number of neurons, is the sigmoid function defined as:
and is the network’s parameter vector of size . For a network of neurons, we have parameters:
where are the output weights, are the neuronal weights, and are the biases of the network.
As our implementation was coded in the Fortran programming language, the Merlin optimization platform [25, 26], which hosts a multitude of minimization algorithms that can be easily combined to form an optimization strategy, was used for training the networks. The solvers that were adopted in our experiments were (i) a pattern search algorithm based on alternating variables, (ii) the irregular simplex method of Nelder and Mead, (iii) Powell’s version of BFGS (TOLMIN) [27] with Goldfarb-Idnani factors and weak Wolfe conditions line search, and (iv) a trust-region version of BFGS based on Powell’ s dogleg technique [28]. All experiments were performed on the PRECIOUS high performance computing infrastructure [29], which consists of 12 computing nodes, each of which utilizes 2 Intel(R) Xeon(R) Gold 5220R processors with 48 computing cores.
Table 1 reports the number of neural form parameters that were used in our experiments. Again, three levels (low, medium, and high) were considered for each test problem.
@@ 6.4 Solution quality criteria
The quality of the obtained trial solution for each test problem was assessed with respect to both residual-related and deviation-related performance measures. Assuming that is the ODE’s residual at , is the number of points over which the performance is evaluated ( for training and for testing), is the exact solution at , is the trial solution at , and , , are the parameters of the neural form, the adopted performance measures are defined as follows:
@@
Residual-related measures
@@
Deviation-related measures
@@
It must be noted that the two measure types assess different performance aspects, as residual-related measures indicate the precision of the ODE’s satisfaction, while deviation-related measures indicate the distance from the exact solution .
@@
| 1 | D | Neural form | 144 | 160 | Ch | 0.11e-14 | 0.19e-13 | 0.13e-15 | 0.19e-13 | ||
| Runge-Kutta | 0.90e-09 | 0.37e-07 | |||||||||
| Neural network | 36 | 160 | Ch | 0.86e-15 | 0.19e-13 | 0.13e-15 | 0.19e-13 | ||||
@@
| 7 | D | Neural form | 288 | 320 | Eq | 0.19e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 | ||
| Runge-Kutta | 0.28e-13 | 0.16e-12 | |||||||||
| Neural network | 288 | 320 | Eq | 0.19e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 | ||||
@@ 7 Experimental results
Our experiments were designed to encompass diverse training scenarios. More specifically, both equidistant grids and Chebyshev grids were used to generate training points. Three levels (low, medium, and high) were considered for the point density as well as for the network size (i.e., its total number of parameters) as reported in Table 1. However, scenarios where the number of network parameters was greater than the number of grid points were not further analyzed, as they are prone to overtraining and, consequently, to declining generalization quality. Thus, 12 different training scenarios were eventually considered.
For each training scenario and test problem, 100 independent experiments were conducted and the obtained solutions were assessed according to the quality measures of Eqs. (34)-(37). The training computational budget of 220000 function evaluations proved to be sufficient for all test problems. Note that Test Problem 2 was also solved using the reductive transformation approach of Section 4, although without significant solution difference.
The results of the proposed neural forms were compared with the corresponding results obtained by (a) the baseline neural method, where the trial solution is a standalone neural network and penalty terms are added to the error function to account for the prescribed conditions, as described by Eq. (8) with , and (b) an explicit fourth-order Runge-Kutta method. Given the large volume of results, Table 2 presents a condensed selection that includes the mean and maximum squared deviations from the exact solutions for the optimal neural form architecture, the corresponding Runge-Kutta method, and the top-performing standalone neural network. The reader can find the detailed results for each test problem in Table 6 of A for the Runge-Kutta method, in Tables 7-20 in B for the baseline neural method (standalone neural network), and in Tables 21-34 of C for the proposed neural forms.
The reported results clearly demonstrate that the proposed optimized neural forms offer effective trial solutions for all test problems. Direct comparisons with the fourth-order Runge-Kutta method offer compelling evidence that neural forms can achieve better accuracy. As we can see in the reported cases in Table 2 as well as in A, neural forms achieved remarkably better deviation from the exact solution on the training dataset in all cases. Another critical advantage of neural forms against Runge-Kutta is that the solutions are provided in closed-formulae, relieving the user from the necessity for further interpolation.
@@
| 1 | D | Neural form | 144 | 160 | Ch | 0.13e-15 | 0.18e-15 | 0.14e-15 | 0.13e-15 | 0.53e-17 | 0.57e-07 | |
| Neural network | 36 | 160 | Ch | 0.13e-15 | 0.27e-10 | 0.48e-12 | 0.14e-15 | 0.29e-11 | ||||
@@
| Min | Neural form | 0.45e-23 | 0.19e-13 | 0.12e-14 | 0.39e-21 | 0.46e-14 | 0.96e-01 |
| MSD | Neural network | 0.13e-22 | 0.17e-08 | 0.12e-09 | 0.19e-21 | 0.41e-09 | |
| Max | Neural form | 0.39e-17 | 0.58e-06 | 0.47e-07 | 0.52e-12 | 0.14e-06 | 0.65e+00 |
| MSD | Neural network | 0.26e-17 | 0.27e-05 | 0.33e-06 | 0.96e-14 | 0.83e-06 | |
| Mean | Neural form | 0.13e-21 | 0.58e-08 | 0.64e-09 | 0.12e-13 | 0.17e-08 | 0.87e+00 |
| MSD | Neural network | 0.58e-19 | 0.68e-06 | 0.50e-07 | 0.95e-15 | 0.16e-06 | |
| Median | Neural form | 0.24-21 | 0.19e-13 | 0.14e-14 | 0.16e-18 | 0.46e-14 | 0.12e+00 |
| MSD | Neural network | 0.19e-20 | 0.55e-06 | 0.39e-07 | 0.49e-19 | 0.13e-06 | |
| St.D. | Neural form | 0.47e-20 | 0.58e-07 | 0.51e-08 | 0.22e-13 | 0.15e-07 | 0.78e+00 |
| MSD | Neural network | 0.35e-18 | 0.53e-06 | 0.55e-07 | 0.15e-16 | 0.14e-06 |
@@
Considering comparisons between neural forms and the baseline neural method using a standalone neural network, a first inspection of Table 2 does not provide conclusive indications of superiority for either of the two methods. More specifically, comparing their MSD values derived from the test datasets suggests that each method achieves better performance for approximately half of the test problems. Further evidence are reported in Table 3, which demonstrates the minimum, maximum, mean, median, and standard deviation of the MSD values over the 100 independent experiments for each test problem and method. The last column reports -values of Wilcoxon rank-sum tests performed between the MSD samples of the compared approaches for each test problem. The small -values obtained in all but two cases indicate statistically significant differences between the two methods for significance level 0.05 and 0.01. In simple words, it appears that for each specific test problem there is a dominant method, although the total successes over the whole test suite are almost equally divided to the two methods.
Taking a step further in our analysis, we examined if there is possible domination of one method over the other on the whole test suite with respect to each specific statistical measure of MSD reported in Table 3. More specifically, each MSD column of Table 3 provides the paired samples of the two neural methods regarding the corresponding statistical measure. For example, the “min” column offers a sample of 16 values for the neural form, and 16 paired values for the neural network. For all these samples, the basic statistical measures were calculated and reported in Table 4. For example, the maximum among the minimum MSD values achieved by the neural form for all test problems was 0.19e-13, while for the basic neural method it was 0.17e-08.
@@
The corresponding paired samples of the two neural methods for each statistical measure were compared using the Wilcoxon paired (signed-rank) test, and the obtained -values are reported in the last column of Table 4. For example, consider the paired samples of the minimum MSD, where each pair consists of the minimum MSD obtained from neural forms and the corresponding minimum MSD of the neural network. Since the calculated -value is 0.96e-1, there is no statistically significant difference in significance level 0.05 or 0.01 between the two methods.
@@
The -values reported in Table 4 for all metrics suggest the lack of significant differences. Therefore, although it appears that for each specific problem it is highly probable to have a winner method, there is no statistically verified dominance of one method over the other across all the test problems. Thus, the proposed neural forms provide solutions of comparable quality to the baseline neural method, while also maintaining the advantage of exact matching of initial/boundary conditions. Note that grids of Chebyshev training points were shown to successfully treat the Runge phenomenon in many cases, thereby enhancing the performance of neural forms with respect to solution deviation metrics.
@@
As previously mentioned, all the reported results were obtained by utilizing the Merlin optimization platform. Notably, widely used solvers such as stochastic gradient descent and the Adam method, which have gained increasing popularity for neural network training, may not achieve good performance. Indeed, Table 35 in D reports the mean and maximum squared deviation from the exact solutions for the best neural form trained using Merlin as described in Section 6.3, and the same neural form trained with Adam. Comparing the reported deviations, it becomes evident that for all test problems Adam either converged to solutions of remarkably lower quality or completely failed to obtain a valid solution.
| Deviation metrics | ||||
| Test Problem | Condition Type | MSD | MXAD | MXAB |
| 1 | Dirichlet | 0.58e-16 | 0.14e-06 | 0.83e-06 |
| 2 | Dirichlet | 0.1087e-20 | 0.1566e-09 | 0.4220e-06 |
| Dirichlet-Neumann | 0.6677e-20 | 0.5358e-09 | 0.6011e-07 | |
| Neumann | 0.8610e-19 | 0.1899e-08 | 0.2850e-05 | |
| Cauchy | 0.3195e-19 | 0.1343e-08 | 0.6592e-08 | |
| Robin | 0.5912e-23 | 0.5281e-11 | 0.1224e-08 | |
@@ 7.1 Deviation bounding
Considered as test cases, we applied the deviation bounding technique presented in Section 5 to the solution of Test Problem 1, and Test Problem 2 for various condition types. Table 5 reports the mean squared deviation (MSD) from the exact solution, as well as two additional metrics. The first metric is the maximum absolute deviation (MXAD) defined as:
and the second metric is the estimated maximum absolute upper bound (MXAB):
where is the exact solution, is the trial solution, and , , are the neural form parameters. For Test Problem 1, the neural form had 36 network parameters, augmented by 18 additional parameters for , while the training dataset consisted of 80 equidistant points in . For Test Problem 2, the neural form had 180 network parameters, augmented by 60 additional parameters for , while the training dataset consisted of 270 equidistant points in . In both cases, the reported metrics MSD, MXAD, and MXAB, were calculated over 1000 equidistant points taken in for Test Problem 1, and in for Test Problem 2. For the latter case, Fig. 2 illustrates the absolute deviation, its calculated bound, and their ratio across the ODE’s domain.
@@ 8 Conclusions
Neural forms constitute an emerging, promising approach for solving ODEs under various condition types. The present work contributes to the existing literature by introducing a systematic procedure for designing proper trial solutions using neural forms, while taking into account the specific type of the prescribed conditions. Also, for boundary value problems, a new technique that transforms a problem with Neumann or mixed boundary conditions into one with parametric Dirichlet boundary conditions was proposed. This transformation provides an alternative path to the proper design of neural forms as trial solutions.
Furthermore, an upper bound for the trial solution’s absolute deviation from the exact one was proposed, providing a new performance measure for the approximation’s quality. For initial value problems, higher accuracy solutions may be achieved by employing domain decomposition techniques. Preliminary results (not reported here) are promising, especially for extended domains, stiff ODEs, and PDEs.
@@
The proposed optimized neural forms have been evaluated on a suite of benchmark tests, consisting of first and second-order ODEs, as well as first-order ODE systems, under a variety of boundary and initial conditions. The obtained solutions were compared to exact ones (wherever available), to solutions obtained by the common penalized neural method (4), and to those obtained by the fourth-order Runge-Kutta algorithm. The comparison clearly confirmed, in agreement with our expectations, the excellent performance of the optimized neural forms. This fact, combined with the closed form type of solution and the exact satisfaction of the boundary or initial conditions, renders this novel approach a highly competitive alternative for solving ODE’s.

### v2

@@ 6.2 Dataset generation
The size of the training set is always problem dependent and, thus, regularly determined through preliminary trial-and-error experimentation. Naturally, complex problems require higher number of training points. Table 1 reports the corresponding numbers in our experiments. For each test problem, three levels (low, medium, and high) were considered for . On the other hand, the test set consisted of equidistant points in all cases.
@@ 6.3 Neural network architecture
where is the number of neurons, is the sigmoid activation function, and is the network’s parameter vector of size . For a network of neurons, we have parameters:
Our implementation was coded in Fortran, while the training was performed by the Merlin optimization platform [26, 27], which offers a variety of robust and efficient minimization routines. The minimization algorithms that were used in our experiments were (i) a pattern search method based on alternating variables, (ii) the irregular simplex method of Nelder and Mead [28], (iii) Powell’s version of BFGS (TOLMIN) [29] with Goldfarb-Idnani factors and weak Wolfe conditions for the line search, and (iv) a trust-region version of BFGS based on Powell’s dogleg technique [30]. All experiments were performed on the PRECIOUS high performance computing infrastructure [31], which consists of 12 computing nodes, each of which utilizes 2 Intel(R) Xeon(R) Gold 5220R processors with 48 computing cores. The number of neural form parameters used in our experiments are reported also in Table 1.
@@ 6.4 Solution quality criteria
An ideal criterion for the quality of an approximate solution is its difference from the exact solution. Given that, in previously unmet problems, the exact solution is not available, an alternative quality criterion can be the size of the ODE residuals. However, this measure may not exhibit the same behavior as the solutions’ difference. In our experiments, all the selected test problems had known exact solutions, hence allowing accurate comparisons and evaluation of our approximate solutions.
Let denote the number of data points over which the performance is evaluated (i.e., for training, and for testing), be the exact solution at , and be the trial solution at , where refers collectively to the neural form parameters. Then, the employed performance measures were as follows:
@@
Measures based on the residuals
@@
Measures based on estimated deviation
(Mean squared estimated deviation) (36) (Maximum squared estimated deviation) (37) -
3.
Measures based on exact deviation
@@
Note that measures based on ODE residuals indicate precision over the ODE’s satisfaction, while measures based on the estimated and exact deviation indicate the distance from the exact solution .
@@
| 1 | D | Neural form | 144 | 160 | Ch | 0.22e-17 | 0.36e-16 | 0.25e-18 | 0.36e-16 | ||
| ode113 | 300 | Ad | 0.58e-28 | 0.18e-26 | 0.12e-23 | 0.99e-21 | |||||
| Neural network | 36 | 80 | Ch | 0.16e-17 | 0.35e-16 | 0.25e-18 | 0.36e-16 | ||||
@@
| Abbreviations: D=Dirichlet, N=Neumann, D-N=Mixed Dirichlet-Neumann, R=Robin, C=Cauchy, Ch=Chebyshev, Eq=Equidistant, Ad=Adaptive | |||||||||||
@@ 7 Experimental results
Our experimental setting was designed to assess the performance of the proposed methodology across a diverse range of problems and computational scenarios that are commonly encountered in practice. More specifically, both equidistant grids and Chebyshev grids were used to generate training points. Three levels (low, medium, and high) were considered for the grid-point density as well as for the network size (total number of parameters) as reported in Table 1. In order to prevent overtraining and its subsequent degradation in generalization quality, scenarios where the number of network parameters exceeded the number of grid points were excluded. Thus, a total number of 12 distinct scenarios were ultimately considered.
For each training scenario and test problem, 100 independent experiments were conducted and the obtained solutions were assessed according to the quality measures presented in Section 6.4. A computational budget of 220000 function evaluations proved to be sufficient for training in all test problems. Note that Test Problem 2 was solved also using the reductive transformation of Section 4, without any significant difference. The results of the proposed neural forms were compared with the corresponding results obtained by:
1.
A baseline neural method where the trial solution is a standalone neural network and penalty terms are added to the error function, in order to account for the prescribed conditions according to the form of Eq. (8) with .
2.
The ode113 numerical solver of Matlab® [23], which is a variable-step, variable-order Adams-Bashforth-Moulton PECE solver of orders 1 to 13. In this case, the boundary value problems were handled via shooting, while the performance over the test datasets was computed using spline interpolation.
Given the large volume of results, Table 2 presents a condensed selection that includes the mean and maximum squared deviations from the exact solutions for the best neural form architecture, the ode113 solver with quintic splines interpolation, and the top-performing standalone neural network. The detailed results for each individual test problem are reported in Table 6 of A for the ode113 solver, in Tables 7-19 of B for the baseline neural method (standalone neural network), and in Tables 20-32 of C for the augmented neural forms.
The reported results clearly demonstrate that the proposed augmented neural forms offered high-quality solutions for all test problems. As expected, the ode113 solver achieved excellent precision at the points of its adaptive grid. However, on the test points where interpolation is necessary, cubic and 7-th order splines yielded inferior results than neural forms. Only quintic splines demonstrated superior performance, as can be verified by inspecting Table 6. It shall be noted that the optimal interpolation choice for this type of numerical methods is not a priori known, hence requiring a trial-and-error experimentation by the user. Also, it shall be underlined that, in contrast to numerical methods, neural forms are interpolants themselves and, thus, they can be differentiated, integrated, or otherwise processed in a straightforward manner. This is an important feature that the traditional numerical solvers lack, thereby emphasizing the complementarity of neural forms to the available numerical solvers.
@@
| 1 | D | Neural form | 144 | 160 | Ch | 0.25e-18 | 0.30e-15 | 0.53e-17 | 0.49e-18 | 0.31e-16 | 0.30e-05 | |
| Neural network | 36 | 80 | Ch | 0.25e-18 | 0.38e-10 | 0.47e-12 | 0.23e-17 | 0.39e-11 | ||||
@@
| Min | Neural form | 0.45e-23 | 0.25e-18 | 0.20e-19 | 0.36e-21 | 0.62e-19 | 0.96e-01 |
| MSD | Neural network | 0.13e-22 | 0.60e-08 | 0.46e-09 | 0.17e-21 | 0.15e-08 | |
| Max | Neural form | 0.39e-17 | 0.58e-06 | 0.50e-07 | 0.10e-11 | 0.15e-06 | 0.68e+00 |
| MSD | Neural network | 0.26e-17 | 0.27e-05 | 0.36e-06 | 0.12e-15 | 0.86e-06 | |
| Mean | Neural form | 0.13e-21 | 0.58e-08 | 0.69e-09 | 0.65e-14 | 0.18e-08 | 0.89e+00 |
| MSD | Neural network | 0.58e-19 | 0.46e-06 | 0.38e-07 | 0.64e-17 | 0.11e-06 | |
| Median | Neural form | 0.24-21 | 0.34e-14 | 0.25e-15 | 0.66e-18 | 0.84e-15 | 0.28e+00 |
| MSD | Neural network | 0.19e-20 | 0.45e-06 | 0.34e-07 | 0.38e-19 | 0.11e-06 | |
| St.D. | Neural form | 0.47e-20 | 0.58e-07 | 0.55e-08 | 0.40e-13 | 0.15e-07 | 0.89e+00 |
| MSD | Neural network | 0.35e-18 | 0.47e-06 | 0.54e-07 | 0.16e-16 | 0.13e-06 |
@@
A comparison of the augmented neural forms against the baseline neural method does not yield a conclusive indication of superiority for either, as can be readily determined by inspecting Table 2. More specifically, the reported MSD values over the test data suggest that each method achieves better performance for approximately half of the test problems. Further evidence is reported in Table 3, which reports the basic descriptive statistics, namely the minimum, maximum, mean, median, and standard deviation of the MSD values over 100 independent experiments, conducted for each test problem and for both the neural forms and the baseline neural method. The last column reports -values of Wilcoxon rank-sum tests performed between the MSD samples of the compared approaches for each test problem. The small -values obtained in all but one case, indicate statistically significant differences between the compared methods for significance levels 0.05 and 0.01. This suggests that, although a dominant method may exist for each distinct test problem, the count over all test problems is almost equally distributed between the two methods.
@@
Further expanding our analysis, we conducted paired (signed-rank) Wilcoxon tests for each one of reported MSD descriptive statistics in Table 3. More specifically, each MSD column of Table 3 provides the paired samples of the two neural methods regarding the corresponding statistic. For example, the “min” column offers a sample of 15 values for the neural form, and 15 paired values for the neural network. For all these samples, the basic statistical measures were calculated and they are reported in Table 4. For example, the maximum among the minimum MSD values achieved by the neural form for all test problems was 0.25e-18, while for the basic neural method it was 0.60e-08. The corresponding paired samples of the two neural methods for each statistic were compared using the Wilcoxon paired (signed-rank) test, and the obtained -values are reported in the last column of Table 4. For example, consider the paired samples of the minimum MSD, where each pair consists of the minimum MSD obtained from neural forms and the corresponding minimum MSD of the neural network. Since the calculated -value is 0.96e-1, there is no statistically significant difference in significance level 0.05 or 0.01 between the two methods.
@@
The -values reported in Table 4 for all metrics suggest the lack of significant differences. Therefore, although it appears that for each specific problem it is highly probable to have a winner method, there is no statistically verified dominance of one method over the other across all the test problems. Thus, the proposed neural forms provide solutions of comparable quality to the baseline neural method, while also maintaining the critical advantage of exact matching the initial/boundary conditions. Note that Chebyshev training grids were shown to successfully treat the Runge phenomenon in many cases, thereby enhancing the performance of neural forms with respect to the solution deviation metrics.
@@
Finally, it is worth noting that widely used solvers such as the stochastic gradient descent, Adam [32], and variations [33], which have recently gained increasing popularity for neural network training in big data applications, may not achieve good performance. Indeed, Table 33 in D reports the mean and maximum squared deviation from the exact solutions for the best neural form trained using Merlin as described in Section 6.3, and the same neural form trained with Adam. Comparing the reported deviations, it becomes evident that for all test problems Adam either converged to solutions of remarkably lower quality or completely failed to obtain a valid solution.
| Deviation metrics | |||||
| Test Problem | Condition Type | MSD | MXD | MSED | MXED |
| 1 | Dirichlet | 0.19e-16 | 0.30e-16 | 0.34e-13 | 0.63e-13 |
| 2 | Dirichlet | 0.11e-20 | 0.24e-19 | 0.62e-15 | 0.26e-14 |
| Dirichlet-Neumann | 0.68e-20 | 0.29e-18 | 0.20e-14 | 0.36e-14 | |
| Neumann | 0.86e-19 | 0.36e-17 | 0.14e-11 | 0.81e-11 | |
| Cauchy | 0.32e-19 | 0.18e-17 | 0.16e-16 | 0.43e-16 | |
| Robin | 0.55e-23 | 0.28e-22 | 0.49e-18 | 0.15e-17 | |
@@ 7.1 Deviation bounding
In order to demonstrate the proposed deviation bounding technique presented in Section 5, we applied it on Test Problems 1 and 2 for various types of conditions. Table 5 reports the mean (MSD), and maximum (MXD) squared deviation from the exact solution, as well as the mean (MSED), and maximum (MXED) estimated deviation. In Test Problem 1, the neural form had 36 network parameters, augmented by 18 additional parameters for , while the training dataset consisted of 80 equidistant points in . In Test Problem 2, the neural form had 180 network parameters, augmented by 60 additional parameters for , while the training dataset consisted of 270 equidistant points in . In both cases, the reported MSD, MXD, MSED, and MXED metrics were calculated over 1000 equidistant points taken in for Test Problem 1, and in for Test Problem 2. For the latter case, Fig. 2 illustrates the absolute deviation and the corresponding estimated upper bound across the ODE’s domain.
In terms of CPU time, both the augmented neural forms and the baseline neural method were quite demanding. However, the baseline method required slightly more processing time due to the need for adjusting the penalty coefficient. The ode113 solver proved to be far more efficient, as expected for methods based on finite differences. However, since finite difference methods do not scale well in many dimensions, the augmented neural forms provide promising alternatives for PDEs.
@@ 8 Conclusions
Augmented neural forms offer a reliable approach for solving ODEs under a variety of boundary and initial conditions. The present work contributes to the existing literature by introducing a systematic procedure for designing appropriate trial solutions with neural forms that exactly satisfy the prescribed problem conditions. Furthermore, a novel technique is presented that transforms a problem with Neumann or Robin boundary conditions into one with parametric Dirichlet conditions. This transformation provides an alternative methodology for formulating proper trial solutions using neural forms. Furthermore, an upper bound for the absolute deviation of the obtained solution from the exact one was introduced, providing a new evaluation metric for the quality of the approximation. For initial value problems, higher accuracy may be achieved by employing domain decomposition techniques.
@@
The proposed augmented neural forms were evaluated on test problems consisting of first- and second-order ODEs, as well as first-order ODE systems, under a variety of boundary and initial conditions. The obtained solutions were compared to the exact ones, as well as to solutions obtained by the common penalized neural method, and to solutions provided by state-of-the-art numerical solvers such as ode113 of Matlab®. The comparison clearly demonstrated that the augmented neural forms can provide accurate solutions with superior generalization performance. Taking also into consideration their closed form type of solution and the exact satisfaction of the boundary/initial conditions, we can infer that the proposed augmented neural forms constitute a robust and reliable alternative for solving ODE’s, while also laying the ground to the treatment of PDEs.